40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 298 of 337  Not logged in ELOG logo
ID Date Author Type Category Subject
  1969   Mon Sep 7 23:18:01 2009 AlbertoUpdateLockingSome locking attempts

Tried to lock the interferometer but arm power didn't get over 65.

Tonight, after the weekend, I resumed the work on locking.

When I started the Mode Cleaner was unlocked because the MZ was also unlocked.

I aligned the MZ and the transmitted power reached about 2.5

Initially the interferometer lost lock at arm power of about 3-4. It looked like the alignment wasn't good enough. So I ran the alignment scripts a few times, first the scripts for the single parts and in the end the one for the full IFO.

Then I also locked again the MZ and this time the transmitted power got to about 4.

In the following locking attempts the the arm power reached 65 but then the lock got lost during the handing of CARM to C1:LSC-PD11_I

I'll keep working on that tomorrow night.

  1968   Mon Sep 7 20:05:18 2009 ranaUpdatePSLRCTEMP v. RMTEMP

Since ~Aug. 27, the reference cavity has been running with no thermal control. This is not really a problem at the 40m; a 1 deg change of the glass cavity

will produce a 5 x 10-7 strain in the arm cavity. That's around 20 microns of length change.

This open loop time gave us the opportunity to see how good our cavity's vacuum can insulation is.

 

rct.pngrct2.png

The first plot below shows the RCTEMP sensors and the RMTEMP sensor. RMTEMP is screwed down to the table close to the can and RCTEMP is on the can, underneath the insulation. I have added a 15 deg offset to RMTEMP so that it would line up with RCTEMP and allow us to see, by eye, what's happening.

There's not enough data here to get a good TF estimate, but if we treat the room temperature as a single frequency (1 / 24 hours) sine wave source, then we can measure the delay and treat it as a phase shift. There's a ~3 hour delay between the RMTEMP and RCTEMP. If the foam acts like a single pole low pass filter, then the phase delay of (3/24)*360 = 45 deg implies a pole at a ~3 hour period. I am not so sure that this is a good foam model, however.

The colorful plot is a scatter plot of RCTEMP v. RMTEMP. The color denotes the time axis - it starts out blue and then becomes red after ten days.

  1967   Fri Sep 4 16:09:26 2009 josephbSummaryVACRebooted RGA computer and reset RGA settings

Steve noticed the RGA was not working today.  It was powered on but no other lights were lit.

Turns out the c0rga machine had not been rebooted when the file system on linux1 was moved to the raid array, and thus no longer had a valid mount to /cvs/cds/.  Thus, the scripts that were run as a cron could not be called.

We rebooted c0rga, and then ran ./RGAset.py to reset all the RGA settings, which had been reset when the RGA had lost power (and thus was the reason for only the power light being lit).

 

Everything seems to be working now.  I'll be adding c0rga to the list of computers to reboot in the wiki.

  1966   Thu Sep 3 23:41:32 2009 AlbertoConfigurationLSCPOX (PD3) aligned

Today I aligned the beam to PD3 (POX) since Steve had moved it.

The DC power read 1.3mV when the beam was on the PD.

  1965   Thu Sep 3 11:20:26 2009 steveUpdateGeneralelectrical ground in place

Quote:

I was told yesterday, that on Friday the construction people accidentally ripped out one of the 40m soil ground.....AND HOW MANY MORE ARE THERE? nobody knows.

It was ~8 ft long and 0.5" diameter buried in the ground. There is no drawing found to identify this exact building ground. They promised to replace this on Wednesday with a 10 ft long and 0.75" diameter.

The  the wall will be resealed where the conduit enters  the  north west corner of the IFO room 104

 

There should be no concern about safety because the 40m building main ground is connected to the CES Mezzanine step-down transformer.

 Atm1 is showing ground bus under N-breaker panel in 40m IFO room north-west corner.

The second ground bus is visible farther down south under M-breaker panel.

Atm2 is the new ground that will be connected to ground bus-N

Attachment 1: Gs-n.JPG
Gs-n.JPG
Attachment 2: newground.JPG
newground.JPG
  1964   Thu Sep 3 10:19:35 2009 steve, albertoUpdatePEMparticle counts and burning hillsides

The San Gabriel mountain has been on fire for  6 days. 144,000 acres of beautiful hillsides burned down and it's still burning.  Where the fires are.

The 40m lab particle counts are more effected by next door building-gardening activity than the fire itself.

This 100 days plot shows that.

Attachment 1: fire6d.jpg
fire6d.jpg
Attachment 2: 40mgarden.JPG
40mgarden.JPG
Attachment 3: 40mgarden2.JPG
40mgarden2.JPG
Attachment 4: RBLOG-FIRE-SKY.JPG
RBLOG-FIRE-SKY.JPG
  1963   Tue Sep 1 13:52:06 2009 steveConfigurationGeneralPOX rfpd is back & needs alignment

Quote:

I removed POX rfpd to see how it is mounted on its base. It is here on the work bench just in case someone wants to use it the IFO over the week end.

 I put POX back to it's place with markers. The pd was removed from it's base so it is for sure misaligned.

  1962   Tue Sep 1 11:23:36 2009 steveUpdateGeneralelectrical ground

I was told yesterday, that on Friday the construction people accidentally ripped out one of the 40m soil ground.....AND HOW MANY MORE ARE THERE? nobody knows.

It was ~8 ft long and 0.5" diameter buried in the ground. There is no drawing found to identify this exact building ground. They promised to replace this on Wednesday with a 10 ft long and 0.75" diameter.

The  the wall will be resealed where the conduit enters  the  north west corner of the IFO room 104

 

There should be no concern about safety because the 40m building main ground is connected to the CES Mezzanine step-down transformer.

Attachment 1: ground.JPG
ground.JPG
  1961   Fri Aug 28 15:30:15 2009 steveConfigurationGeneralPOX rfpd removed

I removed POX rfpd to see how it is mounted on its base. It is here on the work bench just in case someone wants to use it the IFO over the week end.

  1960   Fri Aug 28 13:49:07 2009 robUpdateLockingRF CARM hand off problem

Quote:
Last night, the lock script proceeded to the RF CARM hand-off about half of the time.
However, the hand off was still unsuccessful.

It failed instantly when you turn on the REFL1 input of the CM board, even
when the REFL1 input gain was very low, like -28dB.

I went to the LSC rack and checked the cabling.
The output from the PD11_I (REFL_2) demodulation board is split
into two paths. One goes directly to the ADC and the other one goes
to an SR560. This SR560 is used just as an inverter. Then
the signal goes to the REFL1 input of the CM board.

I found that the SR560 was set to the A-B mode, but B input was open.
This made the signal very noisy. So I changed it to A only mode.
There was also a 1/4 attenuator between the PD11_I output and the SR560.
I took it out and reduced the gain of SR560 from 10 to 2.
These changes allowed me to increase the REFL1 gain to -22dB or so.
But it is still not enough.

I wanted to check the CM open loop TF before the hand-off, but I could
not do that because the lock was lost instantly as soon as I enabled the
test input B of the CM board.
Something is wrong with the board ?

Using the PD11_I signal going into the ADC, I measured the transfer functions
from the CM excitation (digital one) to the REFL_DC (DC CARM signal) and PD11_I.
The TF shapes matched. So the PD11_I signal itself should be fine.

We should try:
* See if flipping the sign of PD11_I signal going to REFL1 input solve the problem.
* Try to measure the CM analog TF again.
* If the noise from the servo analyzer is a problem, try to increase the input gains
of the CM board and reduce the output gain accordingly, so that the signal flowing
inside the CM board is larger.



I'd bet it's in a really twitchy state by the time the script gets to the RF CARM handoff, as the script is not really validated up to that point. It's just the old script with a few haphazard mods, so it needs to be adjusted to accomodate the 15% power drop we've experienced since the last time it was locked.

The CM servo gain needs to be tweaked earlier in the script--you should be able to measure the AO path TF with the arm powers at 30 or so. I was able to do this with the current SR785 setup earlier this week without any trouble.

The 1/4 attenuator is there to prevent saturations on the input to the SR560 when there's still a CARM offset.

Not sure if flipping the sign of PD11 is right, but it's possible we compensated the digital gains and forgot about it. This signal is used for SRCL in the initial acquisition, so we'd have noticed a sign flip.
  1959   Fri Aug 28 12:56:17 2009 YoichiUpdateLockingRF CARM hand off problem
Last night, the lock script proceeded to the RF CARM hand-off about half of the time.
However, the hand off was still unsuccessful.

It failed instantly when you turn on the REFL1 input of the CM board, even
when the REFL1 input gain was very low, like -28dB.

I went to the LSC rack and checked the cabling.
The output from the PD11_I (REFL_2) demodulation board is split
into two paths. One goes directly to the ADC and the other one goes
to an SR560. This SR560 is used just as an inverter. Then
the signal goes to the REFL1 input of the CM board.

I found that the SR560 was set to the A-B mode, but B input was open.
This made the signal very noisy. So I changed it to A only mode.
There was also a 1/4 attenuator between the PD11_I output and the SR560.
I took it out and reduced the gain of SR560 from 10 to 2.
These changes allowed me to increase the REFL1 gain to -22dB or so.
But it is still not enough.

I wanted to check the CM open loop TF before the hand-off, but I could
not do that because the lock was lost instantly as soon as I enabled the
test input B of the CM board.
Something is wrong with the board ?

Using the PD11_I signal going into the ADC, I measured the transfer functions
from the CM excitation (digital one) to the REFL_DC (DC CARM signal) and PD11_I.
The TF shapes matched. So the PD11_I signal itself should be fine.

We should try:
* See if flipping the sign of PD11_I signal going to REFL1 input solve the problem.
* Try to measure the CM analog TF again.
* If the noise from the servo analyzer is a problem, try to increase the input gains
of the CM board and reduce the output gain accordingly, so that the signal flowing
inside the CM board is larger.
  1958   Thu Aug 27 16:14:28 2009 steveSummaryOMCburned photodiode

Old -pre 6/2009  LLO DCPD 3 mm od GTRAN photodiode

Attachment 1: 20090827_173252.jpg
20090827_173252.jpg
Attachment 2: 20090827_170802.jpg
20090827_170802.jpg
  1957   Thu Aug 27 14:00:33 2009 ranaUpdatePSLRC thermal servo impulse response

I stepped the TIDALSET and looked at what happened. Loop was closed with the very low gain.

The RED guy tells us the step/impulse response of the RC can to a step in the heater voltage.

The GREY SLOWDC tells us how much the actual glass spacer of the reference cavity lags the outside can temperature.

Since MINCOMEAS is our error signal, I have upped his SCAN period from 0.5 to 0.1 seconds in the database and reduced its SMOO from 0.9 to 0.0. I've also copied over the Fricke SLOW code and started making a perl PID loop for the reference cavity.

Attachment 1: Untitled.png
Untitled.png
  1956   Thu Aug 27 13:42:08 2009 ranaSummaryPSLReference Cavity Temperature Control: psl.db changes

I made the changes to the psl.db to handle the new Temperature box hardware. The calibrations (EGUF/EGUL) are just copied directly from the LHO .db file (I have rsync'd their entire target area to here).

allegra:c1psl>diff psl.db~ psl.db
341,353d340
< grecord(ai,"C1:PSL-FSS_TIDALOUT")
< {
<       field(DESC,"TIDALOUT- drive to the reference cavity heater")
<       field(DISV,"1")
<         field(SCAN,".5 second")
<       field(DTYP,"VMIVME-3113")
<       field(INP,"#C0 S28 @")
<       field(EGUF,"10")
<       field(EGUL,"-10")
<       field(EGU,"volts")
<       field(LOPR,"-10")
<       field(AOFF,"0")
< }
493,494c480,481
<         field(EGUF,"285.675")
<         field(EGUL,"-214.325")
---
>         field(EGUF,"67.02")
>         field(EGUL,"7.96")
508,509c495,496
<         field(EGUF,"726.85")
<         field(EGUL,"-1273.15")
---
>         field(EGUF,"75.57")
>         field(EGUL,"12.31")
531,532c518,519
<         field(EGUF,"726.85")
<         field(EGUL,"-1273.15")
---
>         field(EGUF,"75.57")
>         field(EGUL,"12.31")
605,617d591
< grecord(ai,"C1:PSL-FSS_TIDALINPUT")
< {
<       field(DESC,"TIDALINPUT- tidal actuator input")
<       field(DISV,"1")
<         field(SCAN,".5 second")
<       field(DTYP,"VMIVME-3123")
<       field(INP,"#C0 S3 @")
<       field(EGUF,"10")
<       field(EGUL,"-10")
<       field(EGU,"volts")
<       field(LOPR,"-10")
<       field(AOFF,"0")
< }
1130a1105,1130
> grecord(ai,"C1:PSL-FSS_TIDALINPUT")
> {
>       field(DESC,"TIDALINPUT- tidal actuator input")
>       field(DISV,"1")
>         field(SCAN,".5 second")
>       field(DTYP,"VMIVME-3123")
>       field(INP,"#C0 S3 @")
>       field(EGUF,"10")
>       field(EGUL,"-10")
>       field(EGU,"volts")
>       field(LOPR,"-10")
>       field(AOFF,"0")
> }
> grecord(ai,"C1:PSL-FSS_TIDALOUT")
> {
>       field(DESC,"TIDALOUT- drive to the reference cavity heater")
>       field(DISV,"1")
>         field(SCAN,".5 second")
>       field(DTYP,"VMIVME-3113")
>       field(INP,"#C0 S28 @")
>       field(EGUF,"10")
>       field(EGUL,"-10")
>       field(EGU,"volts")
>       field(LOPR,"-10")
>       field(AOFF,"0")
> }
1143,1144c1143,1144
<         field(HOPR,"0.010")
<         field(LOPR,"-0.010")
---
>         field(HOPR,"2")
>         field(LOPR,"0")

  1955   Thu Aug 27 12:34:48 2009 YoichiUpdateLockingup to arm power 70
Last night, I tried to run locking scripts.
The power went up to 70 a couple of times .
Then it failed to switch to RF CARM.
I was too tired at that time to figure out what is the problem with the switching.
But it seemed to me that the problem could be solved by some gain tweaking.
Looks like the IFO is back to a good state.
  1954   Wed Aug 26 19:58:14 2009 Rana, AlbertoUpdatePSLReference Cavity Temperature Control: MINCO PID removed

Summary: This afternoon we managed to get the temperature control of the reference cavity working again.

We bypassed the MINCO PID by connecting the temperature box error signal directly into EPICS.

We couldn't configure the PID so that it worked with the modified temperature box so we decided to just avoid using it.

Now the temperature control is done by a software servo by using the channel C1:PSL-FSS_MINCOMEAS as error signal and driving C1:PSL-FSS_TIDALSET (which we have clip-doodle wired directly to the heater input).

 

We 'successfully' used ezcaservo to stabilize the temperature:

ezcaservo -r C1:PSL-FSS_MINCOMEAS -s 26.6 -g -0.00003 C1:PSL-FSS_TIDALSET

 

We also recalibrated the channels:

C1:PSL-FSS_RMTEMP

C1:PSL-FSS_RCTEMP

C1:PSL-FSS_MINCOMEAS

with Peter King on the phone by using ezcawrite (EGUF and EGUL) but we didn't change the database yet. So please do not reboot the PSL computer until we update the database.

 

More details will follow.

Attachment 1: rc.png
rc.png
  1953   Wed Aug 26 16:35:03 2009 AlbertoConfigurationPSLPSL reference cavity temperature box modifications

Basically, in addition to the replacement of the resistors with metal film ones, Peter replaced the chip that provides a voltage reference.

The old one provided about 2.5 V, whereas the new one gets to about 7V. Such reference voltage somehow depends on the room temperature and it is used to generate an error signal for the temperature of the reference cavity.

Peter said that the new higher reference should work better.

  1952   Wed Aug 26 16:31:34 2009 steveUpdatePSLreference cavity temp box temporarly out of order

Quote:

It turned out that half an hour was too long. In less than that the reference cavity temperature passed the critical point when the temperature controller (located just below the ref cav power supply in the same rack) disables the input power to the reference cavity power supply.

The controller's display in the front shows two numbers. The first goes with the temperature of the reference cavity; the second is a threshold set for the first number. The power supply gets enabled only when the first number comes under the threshold value.

Now the cavity is cooling down and it will take about another hour for its temperature to be low enough and for the heater power supply to be powered.

 The cavity temp cooled below SP2 set point 0.1  The Minco SP1 (present temp in Volts) now reading -0.037 so DC power supply was turned on and set to 12V 1A  

 

 

  1951   Wed Aug 26 16:11:41 2009 AlbertoUpdatePSLreference cavity temp box temporarly out of order

It turned out that half an hour was too long. In less than that the reference cavity temperature passed the critical point when the temperature controller (located just below the ref cav power supply in the same rack) disables the input power to the reference cavity power supply.

The controller's display in the front shows two numbers. The first goes with the temperature of the reference cavity; the second is a threshold set for the first number. The power supply gets enabled only when the first number comes under the threshold value.

Now the cavity is cooling down and it will take about another hour for its temperature to be low enough and for the heater power supply to be powered.

  1950   Wed Aug 26 16:10:28 2009 Peter KingConfigurationPSLPSL reference cavity temperature box modifications

The 40m Lab reference cavity temperature box S/N BDL3002 was modified as per DCN D010238-00-C.

These were:

    R1, R2, R5, R6 was 10k now are 25.5k metal film

    R11, R14 was 10k now are 24.9k metal film

    R10, R15 was 10k now are 127k thick film - no metal film resistors available

    R22 was 2.00k now is 2.21k

    R27 was 10k now is 33.2k

    U5, the LM-336/2.5 was removed

    An LT1021-7, 7 V voltage reference was added.  Pin 2 to +15V, pin 4 to ground, pin 6 to U6 pin 3.

    Added an 8.87k metal film resistor between U6 pin 1 and U4 pin 6.

    Added an 8.87k metal film resistor between U6 pin 1 and U4 pin 15.

    The 10k resistor between J8 pin 1 and ground was already added in a previous modification.

In addition R3, R4, R7, R8, R12 and R13 were swapped out for metal film resistors of the same value

(1.00k).

    The jumper connection to the VME setpoint was removed, as per Rana's verbal instructions.

This disables the ability to set the reference cavity vacuum chamber temperature by computer.

 DSC_0731.jpg

 

 

  1949   Wed Aug 26 15:42:17 2009 AlbertoUpdatePSLreference cavity temp box temporarly out of order

Quote:

There's no elog entry about what work has gone on today, but it looks like Peter took apart the reference cavity temperature control around 2PM.

I touched the reference cavity by putting my finger up underneath its sweater and it was nearly too hot to keep my finger in there. I looked at the heater power supply front panel and it seems that it was railed at 30 V and 3 A. The nominal value according to the sticker on the front is 11.5 V and 1 A.

So I turned down the current on the front panel and then switched it off. Otherwise, it would take it a couple of days to cool down once we get the temperature box back in. So for tonight there will definitely be no locking. The original settings are in the attached photo. We should turn this back on with its 1A setting in the morning before Peter starts so that the RC is at a stable temp by the evening. Its important NOT to turn it back on and let it just rail. Use the current limit to set it to 1 A. After the temperature box is back in the current limit can be turned back up to 2A or so. We never need the range for 3A, don't know why anyone set it so high.

While Peter King is still working on the reference cavity temperature box, I turned the power supply for the reference cavity's heater back on. Rana turned it off last night since the ref cav temperature box had been removed.

I just switched it on and turned the current knob in the front panel until current and voltage got back to their values as in Rana's picture.

I plan to leave it like that for half an hour so that the the cavity starts warming up. After that, I'll turn the current back to the nominal value as indicated in the front panel.

  1948   Wed Aug 26 14:45:14 2009 steveUpdatePSLPSL-FSS_RCTEMP of 4 years

The reference cavity vacuum chamber temp is plotted starting Feb 22 of 2005

This plot suggest that the MINCO temp controller is not working properly.

Attachment 1: refcavtemp.jpg
refcavtemp.jpg
  1947   Tue Aug 25 23:16:09 2009 Alberto, ranaConfigurationComputerselog moved in to the cvs path

In nodus, I moved the elog from /export to /cvs/cds/caltech. So now it is in the cvs path instead of a local directory on nodus.

For a while, I'll leave a copy of the old directory containing the logbook subdirectory where it was. If everything works fine, I'll delete that.

I also updated the reboot instructions in the wiki. some of it also is now in the SVN.

  1946   Tue Aug 25 21:55:11 2009 ranaUpdatePSLreference cavity temp box temporarly out of order

There's no elog entry about what work has gone on today, but it looks like Peter took apart the reference cavity temperature control around 2PM.

I touched the reference cavity by putting my finger up underneath its sweater and it was nearly too hot to keep my finger in there. I looked at the heater power supply front panel and it seems that it was railed at 30 V and 3 A. The nominal value according to the sticker on the front is 11.5 V and 1 A.

So I turned down the current on the front panel and then switched it off. Otherwise, it would take it a couple of days to cool down once we get the temperature box back in. So for tonight there will definitely be no locking. The original settings are in the attached photo. We should turn this back on with its 1A setting in the morning before Peter starts so that the RC is at a stable temp by the evening. Its important NOT to turn it back on and let it just rail. Use the current limit to set it to 1 A. After the temperature box is back in the current limit can be turned back up to 2A or so. We never need the range for 3A, don't know why anyone set it so high.

Attachment 1: Untitled.png
Untitled.png
Attachment 2: rc-heater.jpg
rc-heater.jpg
  1945   Tue Aug 25 21:36:28 2009 AlbertoUpdatePSLreference cavity temp box temporarly out of order

 

 Is that the reason of the PSL craziness tonight? See attachment.

Attachment 1: 2009-08-25_PSLtrend.png
2009-08-25_PSLtrend.png
  1944   Tue Aug 25 21:26:12 2009 AlbertoUpdateComputerselog restarted

I just found the elog down and I restarted it.

  1943   Tue Aug 25 18:42:42 2009 steveUpdatePSLreference cavity temp box temporarly out of order

Quote:

The PSL Temperature Box (D980400-B-C, what kind of numbering scheme is that?) modified at LHO/LLO ~8 years ago to have better resolution on the in-loop temperature sensors.

I haven't been able to find a DCN / ECN on this, but there's an elog entry from Hugh Radkins here.  I'm also attaching the PDF of the latest drawing (circa 2000) from the DCC.

The schematic doesn't show it, but I am guessing that the T_SENSE inputs are connected to the AD590 chips, and that 4 of these are attached somehow to the RefCav can. IF this is true, I don't understand why there are input resistors on the LT1125 of U1; the AD590 is supposed to be a current source ?

Peter King is supposed to be coming over to work on this today so whoever spots him should force/cajole/entice him to elog what he's done. Film him if necessary.

 

I also think R1-8 should be swapped into metal film resistors for stability. The datasheet says that it puts out 1 uA/K, so the opamps put out 10 mV/K.

J8 and JP1 should be shorted to disable both the tidal and VME control input. Both are unused and a potential source of drift.

 Peter King is updating our temp box as Hugh did at Hanford Oct.22 of 2001 I still have not seen an updated drawing of this.

The LT 1021-7 reference chip will arrive tomorrow morning. This modification should be completed by noon.

 ** The link to the DCN from Hugh is here in the DCC.

  1942   Tue Aug 25 11:02:39 2009 ranaSummaryPSLTemperature Box

The PSL Temperature Box (D980400-B-C, what kind of numbering scheme is that?) modified at LHO/LLO ~8 years ago to have better resolution on the in-loop temperature sensors.

I haven't been able to find a DCN / ECN on this, but there's an elog entry from Hugh Radkins here.  I'm also attaching the PDF of the latest drawing (circa 2000) from the DCC.

The schematic doesn't show it, but I am guessing that the T_SENSE inputs are connected to the AD590 chips, and that 4 of these are attached somehow to the RefCav can. IF this is true, I don't understand why there are input resistors on the LT1125 of U1; the AD590 is supposed to be a current source ?

Peter King is supposed to be coming over to work on this today so whoever spots him should force/cajole/entice him to elog what he's done. Film him if necessary.

 

I also think R1-8 should be swapped into metal film resistors for stability. The datasheet says that it puts out 1 uA/K, so the opamps put out 10 mV/K.

J8 and JP1 should be shorted to disable both the tidal and VME control input. Both are unused and a potential source of drift.

Attachment 1: D980400-B.pdf
D980400-B.pdf
  1941   Tue Aug 25 03:30:23 2009 YoichiSummaryWIKI-40M UpdateGreen lock and phase noise
While Koji and I were discussing about the green laser lock, we wondered if the common motion of the cavity mirrors,
which won't be suppressed by the green laser servo, will cause any problem to the locking.

Since the common motion of the cavity mirrors is equivalent to the change of the path length from the laser to the
input mirror, it will show up as a phase noise in the error signal.
Unfortunately, since we inject the green laser from the end mirror, this phase noise has opposite sign for the
PSL and the green laser.

I calculated the magnitude of the phase noise using an extremely rough estimate of the common motion of the mirrors.
It is explained in the 40m wiki.
http://lhocds.ligo-wa.caltech.edu:8000/40m/Upgrade_09/GreenLock

The result plot is attached.
(Probably the seismic noise I used is an over estimate.)
Attachment 1: PhaseNoise.png
PhaseNoise.png
  1940   Tue Aug 25 02:37:53 2009 ranaConfigurationComputer Scripts / ProgramsFirefox 3.5 installed for 64 bit linux in apps/
Attachment 1: DSC_0620.JPG
DSC_0620.JPG
  1939   Tue Aug 25 01:27:09 2009 ranaConfigurationComputersRaid update to Framebuilder (not quite)

Quote:

The RAID array servicing the Frame builder was finally switched over to JetStor Sata 16 Bay raid array. Each bay contains a 1 TB drive.  The raid is configured such that 13 TB is available, and the rest is used for fault protection.

The old Fibrenetix FX-606-U4, a 5 bay raid array which only had 1.5 TB space, has been moved over to linux1 and will be used to store /cvs/cds/.

This upgrade provides an increase in look up times from 3-4 days for all channels out to about 30 days.  Final copying of old data occured on August 5th, 2009, and was switched over on that date.

 Sadly, this was only true in theory and we didn't actually check to make sure anything happened.

We are not able to get lookback of more than ~3 days using our new huge disk. Doing a 'du -h' it seems that this is because we have not yet set the framebuilder to keep more than its old amount of frames. Whoever sees Joe or Alex next should ask them to fix us up.

  1938   Tue Aug 25 00:35:04 2009 ranaUpdateGeneralTransfer function of Mode Cleaner Stacks

Looks like all of the accelerometers and seismometers have been disconnected since early AM last Monday when Clara disconnected them for her sensor noise measurement.

Attachment 1: Untitled.png
Untitled.png
  1937   Mon Aug 24 16:48:57 2009 steveHowToVACnew UPS installed

Quote:

As Rob noted last Friday, the UPS which powers the Vacuum rack failed. When we were trying to move the plugs around to debug it, it made a sizzling sound and a pop. Bad smells came out of it.

Ben came over this week and measured the quiescent power consumption. The low power draw level was 11.9 A and during the reboot its 12.2 A. He measured this by ??? (Rob inserts method here).

So what we want is a 120 V * 12.2 A  ~ 1.4 kVA UPS with ~30-50% margin. We look for this on the APC-UPS site:

On Monday, we will order the SUA2200 from APC. It should last for ~25 minutes during an outage. Its $1300. The next step down is $200 cheaper and gives 10 minutes less uptime.

The new APC Smart -UPS 2200VA is now running at  the vacuum rack. There are 2 load monitoring leds on out of 5

Maglev, dry pumps and roughing pumps are not using UPS.

The switch over went smoothly with Yoichi's help.

First we closed all vacuum valves and stopped the two small turbos.

Than turned power off to instruments in the vac-rack and VME: c1vac1 & c1vac2

Maglev was left running.

Now we moved the AC plugs from the wall receptacles over to the back of the UPS and powered them up.

Varian turbos were restarted and vacuum valves were restored in order to reach  vacuum normal condition.

See 40m Vacuum System States and Sequences Manual of 10-24-2001

 

Linux 3 desk top computer is out of order at the pump spool. We should replace it.

The vacuum control screen can be pulled up on a lap top: /cvs/cds/caltech/medm/c0/ve/VacControl_BAK.adj

 

  1936   Mon Aug 24 10:43:27 2009 AlbertoOmnistructureComputersRFM Network Failure

This morning I found that all the front end computers down. A failure of the RFM network drove all the computers down.

I was about to restart them all, but it wasn't necessary. After I power cycled and restarted C1SOSVME all the other computers and RFM network came back to their green status on the MEDM screen. After that I just had to reset and then restart C1SUSVME1/2.

  1935   Fri Aug 21 18:37:16 2009 JenneUpdateGeneralTransfer function of Mode Cleaner Stacks

Using free-swinging Mode Cleaner OSEM data and Guralp seismometers, I have taken transfer functions of the Mode Cleaner stacks.

During this experiment, the MC was unlocked overnight, and one Guralp seismometer was underneath each chamber (MC1/MC3, and MC2).  Clara will let me know what the orientation of the seismometers were (including which seismometer was underneath which chamber and what direction the seismometer axes were pointing), but for now I have included TFs for every combination of suspension motion and seismometer channels.

I combined the 4 OSEM channels for each optic in POS and PIT, and then calibrated each of my sus channels using the method described in Kakeru's elog entry 1413. Units are meters for POS, and radians for PIT.  I also calibrated the guralp channels into meters.

The traces on each plot are: MC_{POS or PIT} / Guralp_{1 or 2}_{direction}.  So each plot shows the coupling between every seismometer direction and a single mirror direction.  The colors are the same for all the plots, ie the gold trace is always Gur1Z.

Attachment 1: TF_osems_guralps.png
TF_osems_guralps.png
  1934   Fri Aug 21 17:49:47 2009 YoichiSummaryComputersUpgrade FE conceptual plan
I started to draw a conceptual diagram of the upgraded FE system.
http://nodus.ligo.caltech.edu:30888/FE/FE-Layout.html
(It takes some time to load the page for the first time)

Some places, tips will pop up when you stop the cursor over an object.
You can also click on the LSC block to see inside.

This is just a start point, so we should add more details.
The source of the diagrams are in the svn:
https://nodus.ligo.caltech.edu:30889/svn/trunk/docs/upgrade08/FE/

I used yEd (http://www.yworks.com/en/products_yed_about.html) to make
the drawings. It is Java, so works on many platforms.
Attachment 1: Picture_2.png
Picture_2.png
  1933   Fri Aug 21 17:28:50 2009 Kevin, ranaSummaryPEMMagnetic Field Measurements Around the Lab

This goal of this test was to measure and map the AC (at 60 Hz) and DC magnetic fields around the interferometer. I've attached the final products which were drawn up with Google SketchUp.

The notes on the maps make them more or less self explanatory: for each numbered point there's an X, Y, and Z measurement produced by the magnetometer. For the AC numbers I measured the Peak-to-Peak value, while for the DC I simply measured the Mean. The magnetometer's axes were always oriented about the same way, with the X arrow on the magnetometer pointing north. I tried to keep variables such as the lights constant as much as possible (they were all on for most measurements, with the exception of a few noted DC ones) and all measurements had the top of the magnetometer at about 32 inches.  The map is pretty close to scale and all the walls and numbered locations were measured out (though the location of objects and the laser tubes is somewhat estimated). I added "landmarks" in the room, which were pretty much the laser tubes, computer racks, and ISC tables.

For each laser room measurement I also took a screenshot using the oscilloscope as a means of recording the shape of the wave for each measurement. Ch1 corresponds to the X value, Ch2 to the Y, and Ch3 to the Z. The screenshots are numbered 1-29 corresponding to the numbers on the map. The zip folders containing the screenshots can be found on the wiki:  PEM:Magnetometers

I should also mention that there is no point #24 and accordingly no 24 screenshot. I realized after I was done that I had messed up the location of that one and instead of risking bad data decided to just remove it.

I decided on the location of the points mainly based on the location of outlets in the room (since I had to plug in the oscilloscope for the AC numbers to set it to 60 Hz). After an initial pass of the room, I went back and filled in some of the larger gaps by moving the magnetometer as far as I could while the oscilloscope remained plugged in to the wall. I used the same points for DC numbers.

Prior to measuring the laser room, I measured the field in other rooms as well. I have

  • AC numbers and screenshots for the control room and the adjoining office room.

  • DC numbers for the entry room and the office room, not including the control room. The X-axis arrow is pointed south (instead of north) for these numbers.

These numbers were sort of a warm up for me to figure out the process and how I would go about recording my data. Since they're not in really important locations and aren't guaranteed to be accurate, I decided not to map them, though the screenshots are still on this Dell Inspiron 1300 Laptop and the measurements in my notebook.

Here are the settings I used on the oscilloscope for all measurements (I merely changed the Vertical Coupling between DC and AC depending on what I was measuring):

  • Impedance: 1M ohms

  • Bandwidth: Full

  • Probe Setup: Voltage 1X

  • Trigger Type: Edge

  • Trigger Coupling: DC

  • Fast Trig: Normal

  • Trigger Mode: Auto

  • Trigger Source: AC Line

  • Acquire Mode: 512 Average

 The notebook that I used contains some additional info that I didn't include in the map, most importantly more precise descriptions of where each of the points is located and the measured distance between each of them (as well as slight changes I made to my measured distances in order to make the room a rectangle; the changes are slight enough that they shouldn't have any real effect on the data).

Since Kevin used our 3-axis Bartington Fluxgate magnetometer, we can guess that we can convert his voltage measurements (below) into magnetic field
by using the manual's guess of 10 uT /V or 10 V/Gauss. This is probably ok at the factor of 2 level, but one day we should calibrate it with a coil.

The punchline is that the DC fields in the lab are roughly what we expect from the Earth's field plus the rebar in our floors: ~1 Gauss. The 60 Hz fields are ~50-500 nT peak-peak.

Attachment 1: AC-field.png
AC-field.png
Attachment 2: DC-field.png
DC-field.png
  1932   Fri Aug 21 17:05:04 2009 JenneUpdateGeneralrestarted the elog

[Kevin, Jenne]

Kevin's awesome final report/elog entry was so awesome that it crashed the elog.  It has been restarted.  We're going to put his pictures and documentation in the wiki, with a link from the elog to prevent re-crashing.

  1931   Thu Aug 20 09:16:32 2009 steveHowToPhotosControl Room Workstation desks lowered to human height

Quote:

There were no injuries...Now we need to get some new chairs.

 The control room desk tops heights on the east side were lowered by 127 mm

 

Attachment 1: P1040788.png
P1040788.png
Attachment 2: P1040782.png
P1040782.png
Attachment 3: P1040786.png
P1040786.png
Attachment 4: P1040789.png
P1040789.png
Attachment 5: P1040785.png
P1040785.png
  1930   Wed Aug 19 23:57:35 2009 robUpdateLockingreport

 

locking work proceeding apace tonight.

diagonalized DRM with setDDphases & senseDRM

initial locks are fairly quick, aqstep script succeeds reliably.

first part of cm_step (handoff CARM-> MCL) usually works.

tuning up later parts of cm_step (presumably due to optical gain changes resulting from MOPA decline). 

got to arm powers ~60.

  1929   Wed Aug 19 18:02:22 2009 JenneUpdateLSCRF PDs aligned

All of the LSC RF PDs have been aligned.  I didn't really change much of anything, since for all of them, the beam was already pretty close to center.  But they all got the treatment of attaching a Voltmeter to the DC out, and adjusting the steering mirror in both pitch and yaw, finding where you fall off the PD in each direction, and then leave the optic in the middle of the two 'edges'.

Before aligning each set (PO, Refl, AS), I followed the procedure in Rob's new RF photodiode Wiki Page

Also, for superstitious reasons, and in case I actually bumped them, I squished all of the ribbon cable connectors into the PDs, just in case.

  1928   Wed Aug 19 17:11:33 2009 JenneUpdateIOOQPDs aligned

Quote:

 

If Rob/Yoichi say the alignment is now good, the we absolutely must center the IOO QPDs and IP POS and IP ANG and MC TRANS  today so that we have good references.

  

 IOO_QPD_POS,    IOO_QPD_ANG,    MC_TRANS,    IP_POS, IP_ANG    have all been centered.

Also, the MCWFS have been centered.

I'm now working on making sure beam is hitting all of the RF PDs around.

  1927   Wed Aug 19 02:17:52 2009 ranaOmnistructureEnvironmentControl Room Workstation desks lowered to human height

There were no injuries...Now we need to get some new chairs.

  1926   Tue Aug 18 19:57:47 2009 rana, JenneUpdatePSLMZ

- we finished the MZ alignment; the contrast is good.

- we did the RFAM tuning using a new technique: a bubble balanced analyzer cube and the StochMon RFPD. This techniques worked well and there's basically no 33 or 166 RFAM. The 133 and 199 are as expected.

- the MC locked right up and then we used the periscope to align to it; the transmission was ~75% of max before periscope tuning. So the beam pointing after the MC should be fine now.

- the Xarm locked up with TRX = 0.97 (no xarm alignment).

 

If Rob/Yoichi say the alignment is now good, the we absolutely must center the IOO QPDs and IP POS and IP ANG and MC TRANS  today so that we have good references.

 

-----------------------------------

The first photo is of our nifty new setup to get the beam to the StochMon PD.  The MZ transmitted beam enters the photo from the bottom right corner, and hits the PBS (which we leveled using a bubble level).  The P-polarization light is transmitted through the cube, and the S-polarization is reflected to the left.  The pure S-polarized light hits a Beam Splitter, which we are using as a pickoff to reduce the amount of light which gets to the PD.  Most of the light is dumped on an aluminum dump.  The remaining light hits a steering mirror (Y1 45-S), goes through a lens, and then hits the StochMon PD.  While aligning the MZ to maximize visibility, we look at the small amount of P-polarized light which passes through the PBS on an IR card, and minimize it (since we want to be sending purely S-polarized light through the EOMs and into the MC).

The second photo is of a spectrum analyzer which is directly connected to the RF out of the StochMon PD.  To minimize the 33MHz and 166MHz peaks, we adjust the waveplates before each of the EOMs, and also adjusted the tilt of the EOM holders.

The final photo is of the EOMs themselves with the Olympus camera.

Once we finished all of our MZ aligning, we noticed that the beam input to the MC wasn't perfect, so Rana adjusted the lower periscope mirror to get the pointing a little better.  

The MZ refl is now at 0.300 when locked.  When Rana reduced the modulation depth, the MZ refl was about 0.050 .  Awesome!

 

Attachment 1: MZ_RFAMmon_setup_small.jpg
MZ_RFAMmon_setup_small.jpg
Attachment 2: MZ_RFAMmon_SpecAnalyzer_small.jpg
MZ_RFAMmon_SpecAnalyzer_small.jpg
Attachment 3: MZ_EOM_IRrefl2_small.jpg
MZ_EOM_IRrefl2_small.jpg
  1925   Tue Aug 18 15:52:27 2009 JenneUpdatePSLMZ
I tweaked up the MZ alignment.  The reflection had been around 0.550, which kept the MEDM indicator green, but was still too high.  I fiddled with BS1, and a little bit with BS2.  When I had the doors of the PSL table open, I got as low as 0.320.  When I closed up and came back to the control room, the MZ refl had drifted up to 0.354.  But it's good again now.

In the future, mirrors shouldn't be so close together that you can't get at their knobs to adjust them No good.  I ended up blocking the beam coming out of the PMC to prevent sticking my hand in some beam, making the adjustment, then removing the dump.  It worked in a safe way, but it was obnoxious. 

  1924   Tue Aug 18 15:16:15 2009 robUpdateIOOMC WFS working again

Rob, Yoichi

 

The MC WFS have apparently been bad for a few days, causing the MC alignment to drift away at DC.  We tried a few things to fix it, including jiggling some EPICS settings in the WFS head & demod screens.  This seemed to work for WFS1 but not WFS2.  Confused, we decided to go stare at the rack 1Y2.  While doing that, we noticed that the top two Sorensens in 1Y1 (these are directly below the Guralp box) were at different voltages from nominal.  The 5V had dropped to 4.2V and the 24V was at 24.6V.  We adjusted the knobs until these were set correctly.  After this, the MC WFS appear to work again.

 

When working in a rack, you must be as careful about accidentally touching things as when working on an optical table.

  1923   Tue Aug 18 14:24:43 2009 YoichiSummaryPSLReference Cavity Inspection
Rana, Koji, Yoichi

To see why the reflected beam from the RC is distorted, we took out
the periscope and the iris in front of the RC. The periscope mirrors
had some gunk and dusts on them. We blew nitrogen air onto them to
remove the dust. Since the gunk did not come off with the air, we
wrapped a Q-tip with lens cleaning paper soaked in Methanol, and wiped
the surface of the mirrors. We did this because it was hard to remove
the mirrors from the periscope (they were in a spring loaded mirror
holders. The springs were too strong to safely remove them without
damaging the mirrors).

Looking into the RC from the front mirror revealed nothing obstructing
in the path.

After the cleaning, we put the periscope back and observed the direct
reflection from the cavity (not locked). It was still distorted
exactly like before.

So we did some tests.
First we injected He-Ne to the RC. It turned out that multiple
reflections from the optical window (not AR coated for He-Ne) made it
almost impossible to investigate anything with He-Ne. But this
observation made us to suspect maybe one side of the window is not AR
coated.

We placed the periscope about 50cm away from the RC and injected the
beam from an angle, so that we can observe the direct reflection.

First, we checked the shape of the beam leaving the periscope. It was good.
We then observed the reflected beam from the RC. It was also good, no distortion.
We made sure that it was really the reflection from the mirror, not from the window
as follows.
We measured the separation between the in coming beam to the cavity and the reflected beam
at two locations. From this, we can guess where the two beams intersect (the reflection point).
The estimated reflection point was far inside the RC enclosure, indicating that it was really
reflection from the front mirror of the RC.
Since we did not see any other reflection beam, we concluded that the AR coating of the window
is good.

We checked the direct reflection beam shape with several different incident angles, but the
beam shape was always good.

We put back the periscope to the original position. This time, we put a high reflectivity mirror
after the output mirror of the periscope. The beam coming out of the circulator (PBS) had a good
circular shape. But still if we let the beam reflected by the cavity, the beam shape is distorted.
Something must be happening in the RC. Unfortunately, we could not figure out what it is.

We put everything back to the original configuration, except for the iris, and the RC alignment
was already good (surprise). After Koji's final tweak, the FSS is now doing fine, but still
the reflected beam is ugly.
  1922   Tue Aug 18 01:16:01 2009 JenneUpdatePSLMach Zehnder is realigned

The Mach Zehnder and I got to know each other today.  The reason for redoing the alignment was to improve pointing from the PSL table into the MC/IFO in hopes that this would solve the MC unlocking problems that we've been having lately.  Since Rana had aligned the IOO QPDs a few weeks ago when all of the alignments and things were good, I used them as a reference for my Mach Zehnder alignment activities. 

 
The order of operations were approximately as follows:

1. Block the secondary (west) arm of the Mach Zehnder using either an aluminum or razor dump.

2. Use SM1 in the MZ to align the beam to the IOO_QPDs (Pos and Ang).  I unfortunately also touched BS2 at this juncture, which made the refl path no longer a reference.

3.  Make sure that the QPD Sum on both Pos and Ang was sensible.  Since there are 2 beamsplitters in a Mach Zehnder, the power on the QPDs should be a quarter when only one beam is on them.  Be careful not to allow the beam no clip on anything.  The biggest problem was the bottom periscope mirror - if you hit it too high or too low, since it is a very thick optic, you end up coming out its side!  This is the frosty part on the edges, totally inappropriate for beams to go through!  Since the side of the periscope mirror isn't HR coated, when going through it like this, I was able to saturate the QPDs.  Not so good. 

4. Also, make sure that this first beam is on the MZ Refl PD.  Do this using the steering optics after the beam has left the MZ.  Use a viewer to look at the PD, and see the small spot of the beam on the diode.  We closed the iris which is present and was standing fully open to remove a spurious beam which was a parallel split-off of the main beam.  Since it was very weak, it is fine.

5.  Unblock the west arm, and block the east arm of the MZ.

6. Align this arm to both the IOO QPDs and the MZ refl diode using the adjustments on BS1, the PZT mirror and if necessary, BS2.  Note that the adjust knobs on the PZT mirror have lock screws.  Make sure to unlock them before adjusting, and relock afterward, to avoid slipping while the PZT is moving.

7.  Unblock all the beams, and make sure there is only one spot both on the transmission side and the reflection side, i.e. the 2 spots from the 2 arms are completely overlapping.  For the Trans side, make sure to look both in the near field and the far field (even after the periscope) to ensure that you really have one spot, instead of just the 2 spots crossing at a single location.

8.  Look at the MZ refl DC out and the PD out from the ISS box (which is essentially MZ trans, looking at Morag and Siobhan) on a 'scope.

9.  Touch / gently wiggle BS1 or another optic, and watch the 'scope.  At the same time, adjust BS1, the PZT mirror and BS2 to maximize the contrast between light and dark fringes.  Ideally, the refl PD should go almost to zero at the dark fringes.

10.  Check that you still have only one overlapping beam everywhere, and that you're actually hitting the MZ refl PD.

11. Because I was concerned about clipping while still figuring out the status of the lower periscope mirror, I removed the beam pipe holders between the last optic before the periscope, and the lower periscope mirror.  The beam pipe had already been removed, this was just the pedestals and the snap-in clamps.

All done for now!  Still to be done:  Optimize the position of the EOMs.  There is a waveplate out front and the EOMs are mounted in such a way that they can be moved in several directions, so that we can optimize the alignment into them.  They ideally only should see a single polarization, in order to apply solely a phase modulation on the beam.  If the input polarization isn't correct, then we'll get a bit of amplitude modulation as well, which on PDs looks like a cavity length change.  Also, the little blue pomona-type box which has the RF signals for the EOMs needs to be clamped to the table with a dog clamp, or better yet needs to be moved underneath the PSL table, with just the cables coming up to the EOMs.  The SMA connections and the SMA cable kept interfering with the MZ refl beam...it's a wonder anyone ever made the beam snake around those cables the way they were in the first place. Right now, the box is sitting just off the side of the table, just inside the doors.

 
Something else that Rana and I did while on the table:  We moved the PMC trans optics just a teensy bit toward the PSL door (to the east) to avoid coming so unbelievably close to the MZ refl optics.  The PMC trans beam shown in the lowest part of my sketch was very nearly clipping on the MZ refl steering optic just near it.  This situation isn't totally ideal, since (as it has been in the past), the first optic which is dedicated to the PMC trans isn't fully sitting on the PSL table.  The pedestal needs to hang off the edge of the table a bit to keep this beam from clipping.  Unfortunately there really isn't space to make a better beam path.  Since we're planning on getting rid of the MZ when the upgrade happens, and this isn't causing us noticeable trouble right now, we're going to let it stay the way it is.

Also, we dumped the reflection from the PMC RFPD onto a razor blade dump. And we noticed that the PZT mirror and BS2 in the MZ are badly vibrationally sensitive. BS2 has a ~400 Hz resonance (which is OK) but a ~150 ms ringdown time!! PZT mirror is similar.

Q = pi * f * tau = 200!  Needs some damping.

Attachment 1: MachZehnderOptics2.pdf
MachZehnderOptics2.pdf
  1921   Mon Aug 17 17:48:49 2009 robSummaryGeneralconlogger restarted

Quote:

I restarted the conlogger on op340m.  This needs to be done when op340m is rebooted--it wasn't done for some reason and so we've lost several days of controls records.

 I added a cronjob on op340m to check every half-hour if the conlog is running, and if not, restart it. 

  1920   Mon Aug 17 17:43:11 2009 ranaSummaryGeneralconlogger restarted

Added the conlog directory to the SVN, minus the enormous data directory. We are now free to make changes to the conlog code.

ELOG V3.1.3-